# Attention-Aided MMSE for OFDM Channel Estimation: Learning Linear Filters with Attention

### v3

@@ IV-C1 MSE and Huber Loss
Possible choices for the loss function $\mathcal{L}(\cdot)$ include MSE and Huber loss [huber1964robust]. We explain these as follows.
  • MSE: For simplicity, we omit the mini batch index $j$ from (32). For a mini-batch of size $\mathcal{B}$, the MSE is defined as
@@
Figure 4: TDL-D: Channel estimation NMSE comparing Huber loss vs. MSE loss
To improve robustness to noise and extreme outliers, and better reflect channel statistics, we train A-MMSE with the Huber loss, which stabilizes learning. As demonstrated in Fig. 4, Huber loss consistently outperform MSE loss, particularly at high SNR where its bounded gradient prevents occasional large residuals from destabilizing the training process-achieving up to $17\%$ lower NMSE at $30$ dB SNR. The detailed training methodology using Huber loss and its implementation for channel estimation are explained in Section VI.
@@ IV-C2 Complexity
The proposed A-MMSE is fundamentally distinguished from existing DNN-based channel estimation approaches. To be specific, the proposed A-MMSE method learns a linear filter matrix $\mathbf{W}_{\text{A-MMSE}}$ during a dedicated training phase. As a result, once trained, channel estimation reduces to a single matrix-vector multiplication, as shown in (19). This enables highly efficient inference, with minimal computational cost and requiring only $2(NM\times L)$ real-valued parameters to store $\mathbf{W}_{\text{A-MMSE}}$. In contrast, as discussed earlier, conventional DNN-based methods [luan2023channelformer, soltani:commlett:19] learn a neural network that maps pilot inputs directly to channel outputs. As a result, each inference (i.e., channel estimation) involves propagating the input through multiple layers of nonlinear operations, leading to higher inference complexity. Specifically, the A-MMSE performs channel estimation using only linear operations with approximately 70,000 $\left(=NM\times L\right)$ parameters, whereas Channelformer requires about 110,000 parameters and includes nonlinear components, such as GELU activations and convolutional layers.
@@ V Rank-Adaptive A-MMSE
In this section, by extending the A-MMSE, we introduce the RA-A-MMSE framework, whose architecture is illustrated in Fig. 5. The RA-A-MMSE incorporates a rank-adaptation (RA) mechanism that adaptively reduces the filter matrix rank according to resource constraints, such as a predefined maximum rank $(r)$ or a limit on the number of filter parameters. By flexibly adjusting the rank to satisfy these constraints, the RA-A-MMSE effectively reduces computational and memory overhead, allowing the filter to operate efficiently even under stringent computation resource limitations. Despite operating at a reduced rank, the RA-A-MMSE still maintains high estimation accuracy. It thus offers a practical and scalable solution for resource-constrained OFDM receivers, achieving a favorable balance between performance and hardware efficiency.
TABLE I: TDL Channel Specification
Scenario Carrier Freq. [Hz] Delay Spread [ns] SCS [Hz] Velocity [km/h] Max Doppler [Hz] Propagation
TDL-A Indoor office 2G 30 15K 5 9.26 NLOS
TDL-B UMi Street-canyon 3.5G 100 30K 30 97.2 NLOS
TDL-C UMa 3.5G 300 30K 100 324.1 NLOS
TDL-D RMa O2I 5G 300 60K 300 1388.9 LOS+NLOS
TDL-E UMa O2I 28G 1000 120K 120 3111.1 LOS+NLOS
@@
To achieve the rank reduction in the RA-A-MMSE, we devise a rank-adaptive (RA) module that produces two trainable matrices, $\mathbf{U}_{r}\in\mathbb{R}^{NM\times r}$ and $\mathbf{V}_{r}\in\mathbb{R}^{L\times r}$, where $r$ denote the desired rank. With $\mathbf{U}_{r}$ and $\mathbf{V}_{r}$, the A-MMSE filter matrix ${\bf{W}}_{\text{A-MMSE}}$ is modified to $\mathbf{W}_{\text{RA-A-MMSE}}=\mathbf{W}_{\text{A-MMSE}}\,{\mathbf{U}_{r}\mathbf{V}_{r}}^{\mathsf{T}}$. Then the rank of $\mathbf{W}_{\text{RA-A-MMSE}}$ is bounded by
@@
Compared to the A-MMSE, the computational cost reduction of the RA-A-MMSE can be quantified by the ratio
$\text{Complexity Cost Ratio:}\;\frac{\text{RA-A-MMSE}}{\text{A-MMSE}}\approx\frac{NMr+Lr}{NML}$ (36)
Considering an OFDM channel comprising $6$ RBs (minimum bandwidth assignment), i.e., $N=72$, $M=14$, and L=36or 72L=36\ $\text${or}\ 72, the RA-A-MMSE with $r=12$ has the reduced computational cost by approximately $35\%$ and $18\%$ for $L=36$ and $L=72$, respectively. Moreover, since the RA module leverages learnable matrices $\left(\mathbf{U}_{r}\ \text{and}\ \mathbf{V}_{r}\right)$ optimized during training, this substantial complexity reduction can be achieved with minimal channel estimation performance loss. We demonstrate this in the next section.
@@
The proposed A-MMSE method exploits sparsity in two primary ways. First, the two-stage Attention encoder implicitly leverages sparsity through its MHA mechanism, which naturally induces soft-sparsity by assigning higher weights to dominant features while suppressing less significant ones. This aligns well with wideband wireless channels that can be captured by a separable, low-rank covariance model ($\mathbf{R}_{\mathrm{full}}\approx\mathbf{R}_{f}\otimes\mathbf{R}_{t}$). Second, the Rank-Adaptive A-MMSE (RA-A-MMSE) framework exploits the rank-deficient structure inherent in the ideal MMSE filter, where all rows share a common whitening filter and differ only by position-dependent correlation coefficients. By learning to identify and operate within this naturally sparse subspace, RA-A-MMSE achieves dramatic rank reduction without significant performance loss. As shown in Section VI, RA-A-MMSE maintains $99.6\%$ of its performance with only $10\%$ of the full rank, demonstrating that it identifies the naturally sparse subspace where MMSE channel estimation resides rather than simply discarding information.
@@ VI Simulation Results
(a) NMSE performance under the TDL-B Scenario (UMi Street-canyon)
(b) NMSE performance under the TDL-D Scenario (RMa O2I)
Figure 7: NMSE comparison over SNR: (a) TDL-B and (b) TDL-D scenarios.
Figure 8: NMSE comparison over SNR: TDL-E scenario.
Figure 9: BER performance under 16-QAM modulation for the 3GPP TDL-E channel model.
@@ VI-A Simulation Setup
For numerical simulations, we follow the 5G NR DM-RS configuration from 3GPP TS 38.211 [3GPP-TS-38.211]. DM-RS’s time-frequency density adapts to channel conditions: front-loaded configurations suit slowly varying channels, while rapidly varying channels require additional symbols to track the dynamics. Frequency selectivity further affects DM-RS pattern design, requiring a trade-off between estimation accuracy and pilot overhead. We illustrate this in Fig. 6.
We evaluate channel estimation performance using TDL channel models (TDL-A to TDL-E) specified in 3GPP TR 38.901, which capture realistic multipath propagation scenarios (see Table I). For each scenario, a continuous sequence of $44,000$ OFDM frames was generated using a standard channel model simulator. By running the channel model continuously without resets and dynamically updating the Doppler shift to model user mobility, we intentionally created a non-stationary channel where the WSSUS assumption is broken globally and holds only locally over short intervals. The generated dataset was partitioned to test generalization: the last 4,000 frames served as the test set with temporal separation from the initial $40,000$ frames, which were divided into $36,000$ training frames and $4,000$ validation frames. This temporal split ensures natural statistical drift between training and test sets, providing a realistic measure of robustness. Model training was conducted using TensorFlow on a single NVIDIA GeForce RTX 4090 GPU.
Due to space limitations, we focus on three representative TDL models: TDL-B (typical urban), TDL-D (high mobility), and TDL-E (mmWave), which together span a broad range of deployment scenarios across frequency bands. The diverse subcarrier spacings specified in Table I, corresponding to different symbol durations, are chosen intentionally to evaluate the proposed method across the fundamental trade-off between Doppler resolution and robustness to channel time-variations. As benchmark methods, we consider the following. For SP-based approaches, we include LS, MMSE, and 1D-MMSE. For DNN-based approaches, we include ChannelNet [soltani:commlett:19] and Channelformer [luan2023channelformer]. Furthermore, in the final subsection, we also compare the performance of a sensing-aided approach [Qing2024IoTJ] and A-MMSE.
(a) NMSE performance under the TDL-B Scenario (UMi Street-canyon) over extended SNR
(b) NMSE performance under the TDL-D Scenario (RMa O2I) over extended SNR
Figure 10: NMSE performance comparison of various channel estimation methods over extended SNR under (a) TDL-B and (b) TDL-D channel scenarios.
(a) NMSE Performance over extended SNR for TDL-E Scenario (mmWave)
(b) Average NMSE vs. Computational Complexity (FLOPs) over all channels, comparing the proposed and conventional estimation curves
Figure 11: NMSE performance over extended SNR (left) and NMSE–complexity trade-off across channels (right) for the proposed and conventional estimators.
VI-B On the Suboptimality of Covariance-Mismatched MMSE Estimation
@@
As described in subsection VI-A, our dataset was intentionally generated to break the WSSUS assumption globally, creating a non-stationary environment where channel statistics vary across OFDM frames. This directly causes covariance mismatch in conventional MMSE estimation, as the filter relies on a fixed estimated covariance $\tilde{\mathbf{R}}$ while the true covariance $\mathbf{R}$ changes over time. The conventional MMSE estimator suffers from an unavoidable performance penalty due to this mismatch, quantified as
$\Delta_{\mathrm{mis}}=\left\|(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\Sigma_{yy}^{1/2}\right\|_{F}^{2}\geq 0.$ (37)
@@
where $\mathbf{W}_{\mathrm{mis}}$ denotes the mismatched filter and $\mathbf{W}^{*}$ represents the oracle MMSE filter with perfect covariance knowledge. As detailed in Appendix B, this penalty is proportional to $\lVert\Delta\mathbf{R}\rVert^{2}_{F}$ and establishes a fundamental performance floor, particularly severe in high-mobility scenarios such as TDL-D and TDL-E.
In contrast, A-MMSE learns the optimal filter directly from data without explicit covariance estimation, bypassing the mismatch problem entirely. The attention mechanism captures true non-stationary channel correlations through an adaptive filter whose error decreases with model capacity and training samples, unlike the fixed penalty $\Delta_{\text{mis}}$ in conventional MMSE. This fundamental distinction explains A-MMSE’s consistent advantage, achieving approximately $43\%$ lower NMSE compared to conventional MMSE across all TDL scenarios.
VI-C Normalized MSE Comparison
For the channel estimation performance comparison, we present the NMSE across different SNR levels in Fig. 7 and Fig. 8. As shown in the figures, the proposed A-MMSE method consistently achieves the best NMSE performance across all SNR regimes and propagation environments, clearly outperforming the existing SP-based approaches as well as the existing DNN-based approaches. Specifically, the A-MMSE achieves around $43\%$ and $48\%$ lower average NMSE compared to the best SP-based baseline (MMSE) and the best DNN-based baseline (Channelformer), respectively. In the considered TDL models, channel statistics vary from slot to slot, rendering classical MMSE no longer strictly optimal. In this situation, the A-MMSE still achieves significantly lower NMSE than MMSE, where the gains are more pronounced in the TDL-D and TDL-E (high-mobility and mmWave). This advantage arises from that the A-MMSE leverages the powerful learning capability from the Attention Transformer. It is also notable that the proposed A-MMSE outperforms Channelformer, which also employs the Attention Transformer architecture. These gains arise from the novel A-MMSE’s two-stage Attention encoder, which first captures frequency-domain correlations and then refines them in the temporal-domain correlation. Crucially, this enables learning the non-stationary channel statistics that vary across OFDM slots more effectively compared to the single-encoder design used in Channelformer.
It is worth noting that the RA-A-MMSE achieves NMSE performance comparable to that of the original A-MMSE, despite its rank reduction. Specifically, in the TDL-B scenario (Fig. 7(a)), the RA-A-MMSE with $10\%$ of rank retains $99.9\%$ of the full-rank A-MMSE, while achieving around $41\%$ lower NMSE than Channelformer. Similar to this, in the TDL-E scenario (Fig. 8) representing mmWave propagation with high frequency selectivity, the RA-A-MMSE with $10\%$ of rank still delivers $99.6\%$ of the full-rank A-MMSE performance and achieves about $25\%$ lower NMSE than Channelformer. These observations highlight the practical advantages of RA-A-MMSE. It preserves the high estimation accuracy of A-MMSE while enabling substantial complexity reduction via rank adaptation.
VI-D BER Performance Analysis
To evaluate the impact of channel estimation accuracy on symbol detection performance, we examine the BER performance for 16-QAM modulation in the TDL-E scenario. Fig. 9 presents the BER curves for the challenging mmWave channel environment. The results reveal distinct performance characteristics at different SNR regimes. At low SNRs (0 - $10$ dB), all methods exhibit similar BER performance, as the noise dominates the overall system performance. However, at high SNRs ($20$ - $30$ dB), where channel estimation accuracy becomes the limiting factor, the A-MMSE demonstrates substantial gains. At $25$ dB, the A-MMSE achieves a BER of $8.59\times 10^{-5}$ compared to $9.89\times 10^{-5}$ for MMSE and $1.17\times 10^{-4}$ for Channelformer, representing improvements of $13.1\%$ and $26.6\%$ respectively. These BER results confirm that the superior channel estimation accuracy of A-MMSE directly translates to enhanced symbol detection performance, with benefits becoming increasingly significant in high SNR.
VI-E Robustness to SNR Mismatch
To evaluate the robustness and generalization of the proposed A-MMSE under SNR mismatch conditions, we consider three independent models trained at SNR values of 0 dB, $15$ dB, and $30$ dB, and evaluate each model across the entire SNR range of 0300-30 dB. Fig. 10 and Fig. 11(a) illustrate the NMSE performance of A-MMSE across the TDL-B, TDL-D, and TDL-E scenarios, respectively.
Across all considered channel profiles, the A-MMSE with SNR mismatch reveals a consistent trend. For instance, the A-MMSE model trained at 0 dB collapses when the test SNR increases. This is mainly because the overwhelming noise obscures the channel’s latent statistics, causing the Attention Transformer to memorize a denoiser rather than learn the true optimal filter. In contrast, the A-MMSE models trained at $15$ dB and $30$ dB generalize effectively to unseen SNR levels. Specifically, in TDL-B, the $15$ dB model retains $99.8\%$ and the $30$ dB model retains $99.3\%$ of the no-mismatch reference performance. In TDL-D, the corresponding performance retentions are $99.1\%$ and $98.1\%$, and in TDL-E, they are $99.0\%$ and $98.1\%$, respectively.
The persistent advantage of the ($15$) dB model stems from a sweet-spot effect, where moderate noise serves as regularization, allowing the network to internalize the full MMSE filter structure without overfitting to noise-free conditions. At ($30$) dB, the regularization effect becomes nearly negligible, resulting in slight robustness loss as operating conditions deviate from the training SNR.
VI-F Performance-Complexity Trade‑off

### v4

@@ IV-C1 MSE and Huber Loss
Possible choices for the loss function $\mathcal{L}(\cdot)$ include MSE and Huber loss [13]. We explain these as follows.
  • MSE: For simplicity, we omit the mini-batch index $j$ from (31). For a mini-batch of size $\mathcal{B}$, the MSE is defined as
@@
To improve robustness to noise and extreme outliers, and better reflect channel statistics, we train A-MMSE with the Huber loss.
@@ IV-C2 Complexity
The proposed A-MMSE is fundamentally distinguished from existing DNN-based channel estimation approaches. To be specific, the proposed A-MMSE method learns a linear filter matrix $\mathbf{W}_{\text{A-MMSE}}$ during a dedicated training phase. As a result, once trained, channel estimation reduces to a single matrix-vector multiplication, as shown in (18). This enables highly efficient inference, with minimal computational cost and requiring only $2(NM\times L)$ real-valued parameters to store $\mathbf{W}_{\text{A-MMSE}}$. In contrast, as discussed earlier, conventional DNN-based methods [25, 36] learn a neural network that maps pilot inputs directly to channel outputs. As a result, each inference (i.e., channel estimation) involves propagating the input through multiple layers of nonlinear operations, leading to higher inference complexity. Specifically, the A-MMSE performs channel estimation using only linear operations with approximately 70,000 $\left(=NM\times L\right)$ parameters, whereas Channelformer requires about 110,000 parameters and includes nonlinear components, such as GELU activations and convolutional layers.
@@ V Rank-Adaptive A-MMSE
In this section, by extending the A-MMSE, we introduce the RA-A-MMSE framework, whose architecture is illustrated in Fig. 4. The RA-A-MMSE incorporates a rank-adaptation mechanism that adaptively reduces the filter matrix rank according to resource constraints, such as a predefined maximum rank $(r)$ or a limit on the number of filter parameters. By flexibly adjusting the rank to satisfy these constraints, the RA-A-MMSE effectively reduces computational and memory overhead, allowing the filter to operate efficiently even under stringent computational resource limitations. Despite operating at a reduced rank, the RA-A-MMSE still maintains high estimation accuracy. It thus offers a practical and scalable solution for resource-constrained OFDM receivers, achieving a favorable balance between performance and hardware efficiency.
TABLE I: Configuration Parameters for HSR and Semi-Urban Scenarios
Scenario Freq. ($f_{c}$) Delay Spread SCS Velocity ($v$) Max Doppler Propagation
HSR 5 GHz 100 ns 60 kHz 350 km/h $\approx 1620.4$ Hz Strong LoS ($K=13$ dB)
Semi-Urban 3.5 GHz 1000 ns 30 kHz 40 km/h $\approx 129.6$ Hz Rich Scattering ($K=3$ dB)
@@
To achieve the rank reduction in the RA-A-MMSE, we devise a rank-adaptive (RA) module that produces two trainable matrices, $\mathbf{U}_{r}\in\mathbb{R}^{NM\times r}$ and $\mathbf{V}_{r}\in\mathbb{R}^{L\times r}$, where $r$ denotes the desired rank. With $\mathbf{U}_{r}$ and $\mathbf{V}_{r}$, the A-MMSE filter matrix ${\bf{W}}_{\text{A-MMSE}}$ is modified to $\mathbf{W}_{\text{RA-A-MMSE}}=\mathbf{W}_{\text{A-MMSE}}\,{\mathbf{U}_{r}\mathbf{V}_{r}}^{\mathsf{T}}$. Then the rank of $\mathbf{W}_{\text{RA-A-MMSE}}$ is bounded by
@@
Compared to the A-MMSE, the computational cost reduction of the RA-A-MMSE can be quantified by the ratio:
$\text{Complexity Cost Ratio:}\;\frac{\text{RA-A-MMSE}}{\text{A-MMSE}}\approx\frac{NMr+Lr}{NML}.$ (35)
Considering an OFDM channel comprising $6$ RBs (minimum bandwidth assignment), i.e., $N=72$, $M=14$, the RA-A-MMSE with $r=12$ reduces the computational cost by approximately $18\%$ for $L=72$. Moreover, since the RA module leverages learnable matrices $\left(\mathbf{U}_{r}\ \text{and}\ \mathbf{V}_{r}\right)$ optimized during training, this substantial complexity reduction can be achieved with minimal channel estimation performance loss. We demonstrate this in the next section.
@@
The proposed A-MMSE method exploits sparsity in two primary ways. First, the two-stage Attention encoder implicitly leverages sparsity through its MHA mechanism, which naturally induces soft-sparsity by assigning higher weights to dominant features while suppressing less significant ones. This aligns well with wideband wireless channels that can be captured by a separable, low-rank covariance model ($\mathbf{R}_{\mathrm{full}}\approx\mathbf{R}_{f}\otimes\mathbf{R}_{t}$). Second, the RA-A-MMSE framework exploits the rank-deficient structure inherent in the ideal LMMSE filter, where all rows share a common whitening filter and differ only by position-dependent correlation coefficients. By learning to identify and operate within this naturally sparse subspace, RA-A-MMSE achieves dramatic rank reduction without significant performance loss. As shown in Section VI, RA-A-MMSE preserves the majority of its performance under reduced-rank operation, demonstrating that it exploits the inherent low-dimensional structure of MMSE channel estimation rather than indiscriminately removing information.
@@ VI Simulation Results
(a) NMSE performance under the Semi-Urban (SU) scenario.
(b) NMSE performance under the High-Speed Rail (HSR) scenario.
Figure 6: NMSE performance comparison of various channel estimation methods over SNR under COST2100 channel models
@@ VI-A Simulation Setup
For numerical simulations, we adopt the 5G NR DM-RS configuration specified in 3GPP TS 38.211 [1]. As illustrated in Fig. 5, the DM-RS time-frequency density is adaptively configured to match channel dynamics, addressing the inherent trade-off between estimation accuracy and pilot overhead. We evaluate channel estimation performance using the geometry-based COST2100 model [22] in Semi-Urban (SU) and High-Speed Rail (HSR) scenarios (Table I). Note that we have also evaluated the proposed method using 3GPP-based TDL channel models. Owing to page limitations, these results are omitted from this paper but are available online [11].
For each scenario, $44,000$ continuous OFDM frames were generated with dynamic Doppler updates. This continuous modeling creates a non-stationary environment where the global WSSUS assumption is explicitly broken, holding only over short local intervals. To assess robustness against statistical drift, we employed a temporal split: the first $40,000$ frames were allocated for training ($36,000$) and validation ($4,000$), while the subsequent $4,000$ frames served as the test set. The model was implemented using TensorFlow on a single NVIDIA RTX 4090 GPU. We note that all the source code are available at [11]. The subcarrier spacings listed in Table I were specifically chosen to evaluate the trade-off between Doppler resolution and robustness to channel time-variations. For benchmarking, we compare our approach against SP-based estimators (LS, Oracle LMMSE, LMMSE mismatch, 1D-LMMSE, and sensing-aided estimation [29]) and DNN-based models (ChannelNet [36] and Channelformer [25]). In particular, the LMMSE mismatch computes the LMMSE filter based on an approximated channel covariance matrix learned from training samples, which results in a mismatch with the true channel statistics.
VI-B Normalized MSE Comparison
For the channel estimation performance comparison, we present the NMSE across different SNR levels in Fig. 6. The results derived from the COST2100 channel model encompass both the SU scenario shown in Fig. 6(a) and the HSR scenario shown in Fig. 6(b). As indicated by the results, the proposed framework achieves lower estimation errors compared to the conventional SP-based and DNN-based baselines across the evaluated channel conditions. Specifically, in the SU scenario, characterized by rich scattering and non-stationary statistics, the A-MMSE achieves an NMSE of $2.07\times 10^{-5}$ at an SNR of $35$ dB. This corresponds to a reduction of approximately $73.8\%$ compared to the mismatched LMMSE estimator ($7.91\times 10^{-5}$) and $82.2\%$ compared to Channelformer ($1.16\times 10^{-4}$).
@@ VI-B Normalized MSE Comparison
The significant performance gains of the proposed method over the mismatched LMMSE stem from the non-stationary nature of the considered channel environment. As described in Subsection VI-A, the WSSUS assumption does not hold globally, and channel statistics vary across OFDM frames. This directly causes covariance mismatch in conventional LMMSE estimation, as the filter relies on a fixed estimated covariance $\tilde{\mathbf{R}}$ while the true covariance $\mathbf{R}$ changes over time. The conventional LMMSE estimator suffers from an unavoidable performance penalty due to this mismatch, quantified as
$\Delta_{\mathrm{mis}}=\left\|(\mathbf{W}_{\mathrm{mis}}-\mathbf{W}^{*})\Sigma_{\mathbf{yy}}^{1/2}\right\|_{F}^{2}\geq 0.$ (36)
@@
where $\mathbf{W}_{\mathrm{mis}}$ denotes the mismatched filter and $\mathbf{W}^{*}$ represents the oracle LMMSE filter with perfect covariance knowledge. This penalty is proportional to $\lVert\Delta\mathbf{R}\rVert^{2}_{F}$ and establishes a fundamental performance floor, which is particularly severe in high-mobility environments such as the COST2100 HSR scenario. In contrast, A-MMSE bypasses the mismatch problem by learning the optimal filter directly from data, effectively capturing non-stationary channel correlations.
Similar to the SU scenario, in the HSR scenario characterized by high-mobility ($350$ km/h), the A-MMSE yields an NMSE of $2.48\times 10^{-4}$ at $35$ dB. This represents a reduction of approximately $39.4\%$ compared to the mismatched LMMSE ($4.09\times 10^{-4}$).
Regarding the RA-A-MMSE, the results indicate that the model retains estimation performance comparable to the full-rank model, even with the reduced filter rank. In the HSR scenario, the RA-A-MMSE with $50\%$ rank achieves an NMSE of $3.01\times 10^{-4}$ at $35$ dB. This corresponds to approximately $82.4\%$ of the estimation accuracy of the full-rank A-MMSE ($2.48\times 10^{-4}$), suggesting that the dominant low-rank structure arising from the strong line-of-sight (LoS) component is effectively captured despite the reduced parameters. Similarly, in the SU scenario, the RA-A-MMSE with $50\%$ rank shows stable estimation accuracy, yielding an NMSE of $4.84\times 10^{-5}$ at $35$ dB. Although the channel exhibits rich scattering, our method maintains an absolute NMSE in the order of $10^{-5}$ (same as original performance) while reducing computational complexity. These observations demonstrate the feasibility of the RA-A-MMSE to balance estimation fidelity and efficiency across different propagation environments.
Remark 5.
We also compare our method against a sensing-aided approach [29], which serves as an advanced SP-based baseline. While this sensing-aided scheme effectively mitigates covariance mismatch—reducing NMSE by approximately $24\%$ (SU) and $12\%$ (HSR) compared to the standard mismatched LMMSE—the proposed A-MMSE significantly outperforms it. Specifically, the A-MMSE achieves an additional NMSE reduction of approximately $65\%$ and $46\%$ in SU and HSR scenarios, respectively, compared to the sensing-aided baseline. This confirms that the data-driven filter learning of A-MMSE offers superior adaptability over purely SP-based refinements.
(a) NMSE performance under the Semi-Urban Scenario over full SNR
(b) NMSE performance under the High-Speed Rail Scenario over full SNR
Figure 7: NMSE performance comparison of various channel estimation methods over extended SNR under COST2100 channel models
VI-C Robustness to SNR Mismatch
To evaluate the robustness of A-MMSE under SNR mismatch, we examine the performance of models trained at fixed SNRs of 0 dB, $20$ dB, and $35$ dB across the full SNR range (Fig. 7). We observe that models trained at extreme SNRs fail to generalize effectively. Specifically, the model trained at 0 dB hits an early error floor in the high-SNR regime, yielding an NMSE of $9.6\times 10^{-3}$ in the SU scenario (Fig. 7(a)) at $35$ dB-approximately $460$ times worse than the matched baseline-because the network prioritizes noise suppression over feature extraction. Conversely, the $35$ dB model suffers in low-SNR conditions due to a lack of noise regularization, exhibiting a $6.7$-fold NMSE degradation in the HSR scenario (Fig. 7(b)) at 0 dB ($0.16$ versus $0.024$).
In contrast, the model trained at $20$ dB achieves a robust trade-off, effectively providing a good operating point. It maintains consistent performance across the entire dynamic range, with only marginal degradation at $35$ dB (e.g., maintaining NMSE within a factor of $1.5$ of the matched baseline in the HSR scenario). We attribute this to the fact that moderate noise serves as regularization, allowing the network to strike a favorable balance between denoising capability and performance in low noise regimes.
We have additionally conducted robustness evaluations under channel-scenario mismatch and developed an online A-MMSE framework to adapt to such conditions. Due to page limitations, these results are provided only online [11].
Figure 8: NMSE–complexity trade-off across channels for the proposed and conventional estimators
VI-D Performance-Complexity Trade‑off